# Full-waveform earthquake source inversion using simulation-based inference

### v1

@@ 5.1 Synthetic Moment Tensor Inversion using Gaussian Noise
To begin with, a simple toy example of moment tensor inversion with fixed source location is explored. First, Gaussian noise is considered; this is particularly appealing since the linearity of the problem with respect to moment tensor components makes inference analytically tractable. In addition to this, optimal score compression saturates the information inequality globally since the first order expansion in Eq. 9 is exact. This means that any departure between the learned posterior and the true posterior is as a result of modelling errors in the NDE . A component of the modelling error should vary as a function of the dataset size , or equivalently the number of simulations.
An artificial pure strike-slip event is constructed with the following moment tensor solution: . Vertical and horizontal component waveforms are simulated with a known Gaussian noise model for a simplified station configuration around the Azores islands, shown in Fig. 5.
The procedure summarised in Fig. 3 is followed, proceeding by uniformly sampling moment tensor components within the hypercube . For each sampled moment tensor , per station per component Gaussian noise is added to the data vector . Fig. 5 shows the full inversion results of this synthetic example, with comparisons between the SBI approach, a long-run MCMC chain with the known Gaussian likelihood, and the analytically known posterior. Both posteriors closely match the true posterior, though the SBI approach does not fully overlap with the analytic result. This minor offset is caused by modelling errors in the neural density estimator (NDE); the NDE has been trained to model the posterior over a very large prior region, so small inaccuracies are unsurprising. It should be noted that while some parameters are not perfectly recovered, this is directly as a result of the noise in the data (a subset of example waveform fits is provided in the Supporting Information, see Fig. SSupporting Information). Noise-free synthetic tests lead to a perfect coincidence between the posterior maximum and the ground truth, as expected.
We then generate independent events within the same prior, and explore how inference performance varies as a function of the number of simulations for each technique. We retrain the same NDE architecture for a range of dataset sizes and perform inference on each event to check the inferred posterior quality. Fig. 6 compares how the sample ensemble of each inversion approach improves with the number of simulations. These results indicate at least an order of magnitude reduction in the number of samples required to generate realistic posterior contours using the SBI approach. Misfits are calculated using the analytically known posterior ], such that , where .
@@ 5.2 Synthetic Moment Tensor Inversion with Real Noise
Fig. 9 indicates substantial issues with the Gaussian likelihood-based approaches. The expected coverage significantly diverges from the ideal distribution at the tails; this indicates that a very high proportion of true model parameters fall well outside the inferred posteriors. This shows that Gaussian likelihood MCMC yields highly overconfident posteriors, which significantly underestimate the uncertainties in the model parameters. This finding is consistent with earlier work which found that an improved covariance parametrisation led to broader posterior uncertainty estimates (Mustać & Tkalčić, 2016). On the other hand, the SBI approach yields very well-calibrated posteriors, with a slight indication of a mild bias, similar to that observed in the Gaussian noise inversions in Fig. 5. Figs. SSupporting Information and SSupporting Information show that Gaussian likelihoods persistently underestimate uncertainties across both land and OBS seismometer arrays, while SBI remains well-calibrated. Fig. SSupporting Information explores the degree of overconfidence with further coverage testing, and demonstrates that the Gaussian likelihood approach appears to overly-constrain each parameter by up to a factor of 3.
@@ 5.3 Real Source Inversions
The non-linear inversion strategy from Section 4.5 is applied to two moderate magnitude events in the region, both of which have GCMT catalogue solutions shown in Fig. 4. The first of these was located slightly north of the Azores islands, while the second occurred just south of Madeira. For both events, all the Azores land stations and UPFLOW OBS (only vertical components to ensure the use of high-quality data) are initially included in the inversion. Observations are then compared with initial synthetics, allowing stations and components to be excluded on the basis of any suspected modelling errors. Note that no data are excluded purely on the basis of a low signal-to-noise ratio, and no windowing is performed, since these effects should not affect the inversion (again, provided they do not obscure any modelling errors).
@@ 5.3.1 North Azores , 13/01/2022
Data from across the seismic array is collected and processed identically as in Sections 5.1 and 5.2 for the event, north of the Azores islands. We ran iterative least squares, using the GCMT solution as the initial time-location, until convergence on a maximum likelihood solution , and constrained with as in Eq. 18. Inference is run using 10,000 simulations for SBI and 160,000 samples with the Gaussian likelihood MCMC approach. Excluding a few minutes to perform iterative least squares, the SBI inversion took minutes ( minutes for dataset generation, minutes training the MAF) and the MCMC approach took minutes. Results are presented in Fig. 10, with the GCMT moment tensor solution presented as reference. We also show a random subselection of data fits comparing the observations with the MLE solution . These show good fits across most of the land station components, as well as a large number of the OBS vertical components across the UPFLOW array.
By virtue of the shared iterative least squares and prior constraining step, both sets of solutions have similar maximum likelihood solutions, but the inferred posteriors diverge significantly. As explored in Section 5.2, the Gaussian likelihood approach yields significantly narrower contours than those produced by the SBI approach. Good compression quality, demonstrated in Fig. SSupporting Information, was found for this event (i.e. the prior constraining step in Eq. 18 ensured low compression loss). This provides strong evidence that the Gaussian likelihood yields highly overconfident posteriors in line with the synthetic inversion experiments carried out in the previous sections. The SBI posterior recovers some of the expected trade-offs, such as between the isotropic components and , maintaining a low volumetric component of the focal mechanism. We observe a wide range of trade-offs that differ between the SBI and Gaussian likelihood approach. These likely reflect genuine features of the posterior caused by the broader contours sampling different regions (and therefore trade-offs) in the domain, and by using an improved empirical likelihood that better captures the effects of seismic noise. Sub-optimal compression will perturb the SBI posterior, likely slightly biasing and blurring it over the domain, but in this instance simple checks ensure this is a minor effect (Fig. SSupporting Information).
@@ 5.3.2 Madeira , 16/02/2022
The iterative least squares procedure was sensitive to the initial conditions, prior, and damping parameters, yielding various least squares solutions converging on false local minima or stationary points. This indicates significant model errors, as well as low sensitivity to some model parameters that made convergence challenging. Once an adequate minimum was found with low misfit values and well-fitting synthetics, poor compression performance was observed, with large compression errors caused by low model sensitivity compounding with non-linearity effects that degrade optimal score compression. This is explored in Fig. SSupporting Information. We therefore selected to mitigate against these problems. We proceeded with inference using 10,000 simulations for SBI and 200,000 samples with the Gaussian likelihood MCMC approach, taking minutes and minutes, respectively. The results are presented in Fig. 11.
Results are broadly similar to the moment tensor inversion in Fig. 10, with the Gaussian likelihood yielding slightly tighter posteriors than the SBI approach. Again, both approaches resolve similar trade-offs between parameters. There are some differences in the posteriors; most notably, the MAP point diverges for some parameters. This deviates from our earlier results, and is caused by the interplay between the non-uniformity introduced by the prior and the empirical model of the likelihood utilised by SBI, . For instance, the SBI MAP longitude shift is closer to the prior mean due to the much broader empirical likelihood the SBI approach produces. This underscores an important point about the advantages of SBI; when complex priors are imposed on the inference problem, a better model of the likelihood can even lead to significantly different MAP solutions in addition to different uncertainty intervals.
@@ 6 Discussion
We have found that across both synthetic and real source inversions, Gaussian-likelihood approaches with various covariance parametrisations are inherently limited for full-waveform moment tensor inversions. Failure to model autocorrelations caused by non-stationary, frequency-specific noise appears to lead to significantly overconfident posteriors. Some prior work has constructed more complex parametrisations of the covariance matrix by fitting tapered sinusoids to the autocorrelation structure (Mustać & Tkalčić, 2016). However, a more general approach to empirically modelling a Gaussian covariance is challenging due to the practical difficulties in estimating Gaussian auto-covariances of time series from limited samples (Wu & Xiao, 2012). Both these approaches may be hamstrung by the strong non-stationarity of the frequency content and amplitude of seismic noise. These factors suggest that wholesale empirical modelling of the likelihood, as performed by SBI, is highly desirable in the full-waveform inversion context. Additionally, further work could explore how these results vary under different frequency ranges. For instance, shorter period seismic waveforms close to the microseismic noise band may be noisier but more stationary than for the wave periods considered in this study (Peterson, 1993). This could potentially lead to differences in the performance of standard Gaussian likelihood-based techniques.
@@ 6.1.1 Optimal Score Compression
These issues are inherent limitations of optimal score compression when applied to non-linearities in the observations. Often, practitioners perform two-step compression where the raw observations are transformed into a first set of more stable summaries, after which the linearising assumption in optimal score compression is adequate (Alsing et al., 2019). This approach may reduce the discriminative power of SBI but ensures its robustness. In seismology, this may involve utilising wavelet transforms and power spectra in conjunction with classical compression techniques such as principal component analysis (Vavryčuk et al., 2017; Yu et al., 2018; Piras et al., 2023). Another approach could aim to adapt compression techniques to more time-shift robust techniques that utilise the commonly used decorrelation rather than distance (Stähler & Sigloch, 2014; Phạm, 2024) or perhaps take inspiration from recent work using optimal transport in seismology (Engquist et al., 2016; Sambridge et al., 2022). Alternatively, a more general approach could utilise end-to-end deep learning architectures inspired by e.g. Münchmeyer et al. (2021); Nooshiri et al. (2022) to directly learn how to compress these high-dimensional seismic observations using ML (Jeffrey et al., 2020).
@@ 6.1.2 Addressing Modelling Errors
One key limitation of this study is that it only partially addressed issues related with modelling errors (e.g., inaccurate earth structure used, missing crucial physical processes such as seismic wave propagation in the oceans, etc). Modelling errors can cause serious inaccuracies in seismic source inversion, potentially leading to highly overconfident, inaccurate posteriors. A wide range of practical efforts to address this issue have been utilised across the field. Valentine & Trampert (2012) explored ensembles of “plausible” solutions; Duputel et al. (2012, 2014); Vasyura-Bathke et al. (2021); Phạm et al. (2024) incorporated an additional “corrective” Gaussian terms in the likelihood that addresses auto-correlated errors arising from modelling errors (e.g. errors caused by phase-offsets), and demonstrate significantly improved robustness to model errors. These approaches, while essential to effectively treat moment tensor inversions under earth structure errors, involve simplifications and assumptions about the effect of model errors on the inference process. They could therefore produce uncertainty intervals that are difficult to interpret. The approach presented by Stähler & Sigloch (2014, 2016) remains robust and principled in dealing with modelling errors in the Bayesian framework, as they demonstrate that modelling the distribution over decorrelation errors between observed seismograms and synthetics prove more robust than Gaussian likelihood noise models. The success of this approach, however, hinges on the availability of an accurate, curated catalogue, which may be infeasible to acquire or whose biases may be propagated directly into the source parameter estimates.
@@ Acknowledgements.
We thank the UPFLOW project, funded by the European Research Council under the European Union’s Horizon 2020 research and innovation program (grant agreement No 101001601). AAS was supported by the STFC UCL Centre for Doctoral Training in Data Intensive Science. DP was supported by a Swiss National Science Foundation (SNSF) Professorship grant (No. 202671), and by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”.References
@@
Figure S4: Examples of artificial noisy observations, one with known Gaussian noise and another using real seismic noise data, presented alongside the corresponding best waveform fits.

### v2

@@ 5.1 Synthetic Moment Tensor Inversion using Gaussian Noise
To begin with, a simple toy example of moment tensor inversion with fixed source location is explored. First, Gaussian noise is considered; this is particularly appealing since the linearity of the problem with respect to moment tensor components makes inference analytically tractable (Tarantola, 2005). In addition to this, optimal score compression saturates the information inequality globally since the first order expansion in Eq. 10 is exact. This means that any departure between the learned posterior and the true posterior is as a result of modelling errors in the NDE . A component of the modelling error should vary as a function of the dataset size , or equivalently the number of simulations.
An artificial pure strike-slip event is constructed with the following moment tensor solution: . Vertical and horizontal component waveforms are simulated with a known Gaussian noise model for a simplified 13 station configuration around the Azores islands, shown in Fig. 5.
The procedure summarised in Fig. 3 is followed, proceeding by uniformly sampling moment tensor components within the hypercube . For each sampled moment tensor , per station per component Gaussian noise is added to the data vector . Fig. 5 shows the full inversion results of this synthetic example, with comparisons between the SBI approach, a long-run MCMC chain with the known Gaussian likelihood, and the analytically known posterior. Both posteriors closely match the true posterior, though the SBI approach does not fully overlap with the analytic result. This minor offset is caused by modelling errors in the neural density estimator (NDE); the NDE has been trained to model the posterior over a very large prior region, so small inaccuracies are unsurprising. It should be noted that while some parameters are not perfectly recovered, this is directly as a result of the noise in the data. Noise-free synthetic tests lead to a perfect coincidence between the posterior maximum and the ground truth, as expected.
We then generate independent events within the same prior, and explore how inference performance varies as a function of the number of simulations for each technique. We retrain the same NDE architecture for a range of dataset sizes and perform inference on each event to check the inferred posterior quality. Fig. 6 compares how the sample ensemble of each inversion approach improves with the number of simulations. These results indicate at least an order of magnitude reduction in the number of samples required to generate realistic posterior contours using the SBI approach. Misfits are calculated using the analytically known posterior ], such that , where (Tarantola, 2005).
@@ 5.2 Synthetic Moment Tensor Inversion with Real Noise
Fig. 9 indicates substantial issues with the Gaussian likelihood-based approaches. The expected coverage significantly diverges from the ideal distribution at the tails; this indicates that a very high proportion of true model parameters fall well outside the inferred posteriors. This shows that Gaussian likelihood MCMC yields highly overconfident posteriors, which significantly underestimate the uncertainties in the model parameters. This finding is consistent with earlier work which found that an improved covariance parametrisation led to broader posterior uncertainty estimates (Mustać & Tkalčić, 2016). On the other hand, the SBI approach yields very well-calibrated posteriors, with a slight indication of a mild bias, similar to that observed in the Gaussian noise inversions in Fig. 5. Figs SSupporting Information and SSupporting Information show that Gaussian likelihoods persistently underestimate uncertainties across both land and OBS seismometer arrays, while SBI remains well-calibrated. Fig. SSupporting Information explores the degree of overconfidence with further coverage testing, and demonstrates that the Gaussian likelihood approach appears to overly-constrain each parameter by up to a factor of 3.
@@ 5.3 Real Source Inversions
The non-linear inversion strategy from Section 4.5 is applied to two moderate magnitude events in the region, both of which have GCMT catalogue solutions shown in Fig. 4. The first of these was located slightly north of the Azores islands, while the second occurred just south of Madeira. For both events, all the Azores land stations and UPFLOW OBS (only vertical components to ensure the use of high-quality data) are initially included in the inversion. Observations are then compared with initial synthetics, allowing stations and components to be excluded on the basis of any clear modelling errors by visual inspection. Note that no data are excluded purely on the basis of a low signal-to-noise ratio, and no windowing is performed, since these effects should not affect the inversion (again, provided they do not obscure any modelling errors). The resulting subset of accepted stations and components is then fixed and used throughout all dataset generation and inference.
@@ 5.3.1 North Azores , 13/01/2022
Data from across the seismic array is collected and processed identically as in Sections 5.1 and 5.2 for the event, north of the Azores islands. We ran iterative least squares, using the GCMT solution as the initial time-location, until convergence on a maximum likelihood solution , and constrained with as in Eq. 18. Inference is run using 10,000 simulations for SBI and 160,000 samples with the Gaussian likelihood MCMC approach. Excluding a few minutes to perform iterative least squares, the SBI inversion took minutes ( minutes for dataset generation, minutes training the MAF) and the MCMC approach took minutes. Results are presented in Fig. 10, with the GCMT moment tensor solution presented as reference (note that GCMT inversions constrain the volumetric component to zero). We also show a random subselection of data fits comparing the observations with the MLE solution . These show good fits across most of the land station components, as well as a large number of the OBS vertical components across the UPFLOW array.
By virtue of the shared iterative least squares and prior constraining step, both sets of solutions have similar maximum likelihood solutions, but the inferred posteriors diverge significantly. As explored in Section 5.2, the Gaussian likelihood approach yields significantly narrower contours than those produced by the SBI approach. Good compression quality, demonstrated in Fig. SSupporting Information, was found for this event (i.e. the prior constraining step in Eq. 18 ensured low compression loss). This provides strong evidence that the Gaussian likelihood yields highly overconfident posteriors in line with the synthetic inversion experiments carried out in the previous sections. The SBI posterior recovers some of the expected trade-offs, such as between the isotropic components and . We find a large volumetric component across both ensembles; it is unclear to what degree this is a genuine feature of the event, or whether it is driven by insufficient data and modelling errors. Future work could investigate solution sensitivity to volumetric component by enforcing traceless priors during the inversions. Further interpretation, such as attributing the focal mechanism to tensional crack, is challenging without constraining information on subsurface structures, since the classical moment tensor model of Aki & Richards (2002) may not be applicable (Tape & Tape, 2013; Alvizuri & Tape, 2016). We observe a wide range of trade-offs that differ between the SBI and Gaussian likelihood approach. These likely reflect genuine features of the posterior caused by the broader contours sampling different regions (and therefore trade-offs) in the domain, and by using an improved empirical likelihood that better captures the effects of seismic noise. Sub-optimal compression will perturb the SBI posterior, likely slightly biasing and blurring it over the domain, but in this instance simple checks ensure this is a minor effect (Fig. SSupporting Information).
@@ 5.3.2 Madeira , 16/02/2022
The iterative least squares procedure was sensitive to the initial conditions, prior, and damping parameters, yielding various least squares solutions converging on false local minima or stationary points. This indicates significant model errors, as well as low sensitivity to some model parameters that made convergence challenging. Once an adequate minimum was found with low misfit values and well-fitting synthetics, poor compression performance was observed, with large compression errors caused by low model sensitivity compounding with non-linearity effects that degrade optimal score compression. This is explored in Fig. SSupporting Information. We therefore selected to mitigate against these problems. We proceeded with inference using 10,000 simulations for SBI and 200,000 samples with the Gaussian likelihood MCMC approach. Inference took minutes (including dataset generation, training, and sampling) and minutes, respectively. The results are presented in Fig. 11.
Results are broadly similar to the moment tensor inversion in Fig. 10, with the Gaussian likelihood yielding slightly tighter posteriors than the SBI approach. Both sets of solutions show little variability, driven by high signal-to-noise ratios and poor sensitivities for this event; the lune plot in Fig. 11 indicates the two approaches yield very similar mechanisms. Again, both approaches resolve similar trade-offs between parameters. There are some differences in the posteriors; most notably, the MAP point is different for some parameters. This deviates from our earlier results, and is caused by the interplay between the non-uniformity introduced by the prior and the empirical model of the likelihood utilised by SBI, . For instance, the SBI MAP longitude shift is closer to the prior mean due to the much broader empirical likelihood the SBI approach produces. This underscores an important point about the advantages of SBI; when complex priors are imposed on the inference problem, a better model of the likelihood can even lead to significantly different MAP solutions in addition to different uncertainty intervals.
@@ 6 Discussion
We have found that across both synthetic and real source inversions, Gaussian-likelihood approaches with various covariance parametrisations are inherently limited for full-waveform moment tensor inversions. Failure to model autocorrelations caused by non-stationary, frequency-specific noise appears to lead to significantly overconfident posteriors. Some prior work has constructed more complex parametrisations of the covariance matrix by fitting tapered sinusoids to the autocorrelation structure (Mustać & Tkalčić, 2016). More generic empirical modelling of the Gaussian covariance is challenging due to the practical difficulties in estimating Gaussian auto-covariances of time series from limited samples (Wu & Xiao, 2012). Transdimensional inference, which jointly inverts the noise and model parameters hierarchically, could also be explored for seismic source inversion for better estimates of the noise model (Sambridge et al., 2006, 2013). However, all these approaches may be hamstrung by the strong non-stationarity of the frequency content and amplitude of seismic noise, as well as potential inherent non-Gaussianity. These factors suggest that wholesale empirical modelling of the likelihood, as performed by SBI, is highly desirable in the full-waveform inversion context. Additionally, further work could explore how these results vary under different frequency ranges. For instance, shorter period seismic waveforms close to the microseismic noise band may be noisier but more stationary than for the wave periods considered in this study (Peterson, 1993). This could potentially lead to differences in the performance of standard Gaussian likelihood-based techniques.
@@ 6.1.1 Optimal Score Compression
These issues are inherent limitations of optimal score compression when applied to non-linearities in the observations. Often, practitioners perform two-step compression where the raw observations are transformed into a first set of more stable summaries, after which the linearising assumption in optimal score compression is adequate (Alsing et al., 2019). This approach may reduce the discriminative power of SBI but ensures its robustness. In seismology, this may involve utilising wavelet transforms and power spectra in conjunction with classical compression techniques such as principal component analysis (Vavryčuk et al., 2017; Yu et al., 2018; Piras et al., 2023). Another approach could aim to adapt compression techniques to more time-shift robust techniques that utilise the commonly used decorrelation rather than distance (Stähler & Sigloch, 2014; Phạm, 2024) or perhaps take inspiration from recent work using optimal transport in seismology (Engquist et al., 2016; Sambridge et al., 2022). Alternatively, a more general approach could utilise end-to-end deep learning architectures inspired by e.g. Münchmeyer et al. (2021); Nooshiri et al. (2022) to directly learn how to compress these high-dimensional seismic observations using ML (Jeffrey et al., 2020). Such deep learning approaches could be integrated into an SBI framework to make inference robust to data artefacts and glitches, mitigating against the need for manual data inspection and selection (as in e.g., McBrearty & Beroza, 2022, 2023).
@@ 6.1.2 Addressing Modelling Errors
One key limitation of this study is that it only partially addressed issues related with modelling errors (e.g., inaccurate earth structure used, missing crucial physical processes such as seismic wave propagation in the oceans, etc). Modelling errors can cause serious inaccuracies in seismic source inversion, potentially leading to highly overconfident, inaccurate posteriors. A wide range of practical efforts to address this issue have been utilised across the field. Valentine & Trampert (2012) explored ensembles of “plausible” solutions; Duputel et al. (2012, 2014); Vasyura-Bathke et al. (2021); Phạm et al. (2024) incorporated an additional “corrective” Gaussian terms in the likelihood that addresses auto-correlated errors arising from modelling errors (e.g. errors caused by phase-offsets), and demonstrate significantly improved robustness to model errors. However, since for instance waveform time shifts do not comply with a Gaussian distribution (Yagi & Fukahata, 2008; Hu et al., 2023), these approaches represent simplifications and assumptions about the effect of model errors on the inference process. They could therefore produce incorrect uncertainty intervals that are difficult to interpret. The approach presented by Stähler & Sigloch (2014, 2016) remains robust and principled in dealing with modelling errors in the Bayesian framework, as they demonstrate that modelling the distribution over decorrelation errors between observed seismograms and synthetics prove more robust than Gaussian likelihood noise models. The success of this approach, however, hinges on the availability of an accurate, curated catalogue, which may be infeasible to acquire or whose biases may be propagated directly into the source parameter estimates.
@@ Acknowledgements.
We thank the UPFLOW project, funded by the European Research Council under the European Union’s Horizon 2020 research and innovation program (grant agreement No 101001601). AAS was supported by the STFC UCL Centre for Doctoral Training in Data Intensive Science (grant ST/W00674X/1) and by departmental and industry contributions. AAS was also supported by the A. G. Leventis Foundation educational grant scheme. DP was supported by a Swiss National Science Foundation (SNSF) Professorship grant (No. 202671), and by the SNF Sinergia grant CRSII5-193826 “AstroSignals: A New Window on the Universe, with the New Generation of Large Radio-Astronomy Facilities”.References
@@
- Alvizuri & Tape (2016) Alvizuri, C. & Tape, C., 2016. Full moment tensors for small events (m w¡ 3) at uturuncu volcano, bolivia, Geophysical Journal International, 206(3), 1761–1783.
@@
- Cabieces et al. (2024) Cabieces, R., Harris, K., Ferreira, A., Tsekhmistrenko, M., Hicks, S., Krüger, F., Geissler, W., Hannemann, K., & Schmidt-Aursch, M., 2024. Clock drift corrections for large aperture ocean bottom seismometer arrays: application to the upflow array in the mid-atlantic ocean, Geophysical Journal International, 239(3), 1709–1728.
@@
- Fluri et al. (2021) Fluri, J., Kacprzak, T., Refregier, A., Lucchi, A., & Hofmann, T., 2021. Cosmological parameter estimation and inference using deep summaries, Physical Review D, 104(12), 123526.
@@
- Giardini (1984) Giardini, D., 1984. Systematic analysis of deep seismicity: 200 centroid-moment tensor solutions for earthquakes between 1977 and 1980, Geophysical Journal International, 77(3), 883–914.
@@
- Lu et al. (2022) Lu, T., Haiman, Z., & Zorrilla Matilla, J. M., 2022. Simultaneously constraining cosmology and baryonic physics via deep learning from weak lensing, Monthly Notices of the Royal Astronomical Society, 511(1), 1518–1528.
@@
- McBrearty & Beroza (2022) McBrearty, I. W. & Beroza, G. C., 2022. Earthquake location and magnitude estimation with graph neural networks, in 2022 IEEE international conference on image processing (ICIP), pp. 3858–3862, IEEE.
- McBrearty & Beroza (2023) McBrearty, I. W. & Beroza, G. C., 2023. Earthquake phase association with graph neural networks, Bulletin of the Seismological Society of America, 113(2), 524–547.
@@
- Sambridge et al. (2006) Sambridge, M., Gallagher, K., Jackson, A., & Rickwood, P., 2006. Trans-dimensional inverse problems, model comparison and the evidence, Geophysical Journal International, 167(2), 528–542.
- Sambridge et al. (2013) Sambridge, M., Bodin, T., Gallagher, K., & Tkalčić, H., 2013. Transdimensional inference in the geosciences, Philosophical Transactions of the Royal Society A: Mathematical, Physical and Engineering Sciences, 371(1984), 20110547.
@@
- Sharma et al. (2024) Sharma, D., Dai, B., & Seljak, U., 2024. A comparative study of cosmological constraints from weak lensing using convolutional neural networks, Journal of Cosmology and Astroparticle Physics, 2024(08), 010.
@@
- Tape & Tape (2012) Tape, W. & Tape, C., 2012. A geometric setting for moment tensors, Geophysical Journal International, 190(1), 476–498.
- Tape & Tape (2013) Tape, W. & Tape, C., 2013. The classical model for moment tensors, Geophysical Journal International, 195(3), 1701–1720.
@@
- Yagi & Fukahata (2008) Yagi, Y. & Fukahata, Y., 2008. Importance of covariance components in inversion analyses of densely sampled observed data: an application to waveform data inversion for seismic source processes, Geophysical Journal International, 175(1), 215–221.